A new car is purchased for 24800 dollars. The value of the car depreciates at 12% per year. What will the value of the car be, to the nearest cent, after 13 years?


1
Expert's answer
2021-04-14T14:48:11-0400

An=A0(1 - p/100)n

An=24800(1 - 0.12)13=4706.81 dollars